OpenStudy (anonymous):

If a + b = 9 and b = 3, then 4a =

OpenStudy (solomonzelman):

plug in 3 instead of b, into `a+b=9` solve for a. multiply the result times 4, to get the value of `4a`

OpenStudy (anonymous):

which number i should multiply?

OpenStudy (solomonzelman):

can you do step 1? what do you get after you do step 1 ?

OpenStudy (anonymous):

i really mixed it up im sory i need help!

OpenStudy (solomonzelman):

\(\Large\color{blue}{ a + b = 9 }\) you are given that \(\Large\color{blue}{ b = 3 }\) So, •••• \(\Large\color{blue}{ a + b = 9 }\) \(\Large\color{blue}{ a + \color{red}{ 3 } = 9 }\) (plugged in 3 for x) \(\Large\color{blue}{ a + 3 = 9 }\) \(\Large\color{blue}{ a + 3 \color{red}{ - 3 } = 9 \color{red}{ -3 } }\) \(\Large\color{blue}{ a = 6}\)

OpenStudy (solomonzelman):

So if `a` is 6, then 4a= ?

OpenStudy (anonymous):

a=24

OpenStudy (anonymous):

thanks

geerky42 (geerky42):

4a = 24*

OpenStudy (anonymous):

yap correct

OpenStudy (solomonzelman):

yes, correct @JRDetroit 4a=24 :)

OpenStudy (anonymous):

how long you will be online?

OpenStudy (anonymous):

If f(x) = 3x3 – 3x, then f(2) equals?

OpenStudy (anonymous):

anyone can solve this?

OpenStudy (goformit100):

Sir/Ma'am, A Warm Welcome To Open Study.